What mass of neon gas is required to fill a 4.32-L container to a pressure of 1.36 atm at 21°C?
Answer #1

ne PV RT (1.36 atm) (4.32L) (0.08 21 L-atmimod! Kt) (21+273) k n= 0-24 34 moles mass= nx atomic mass of He = 0.2434 modx 20.1According to ideal gas law PV = nRT p is the pressure V is the volume n is number of moles R = 0.0821 L. atm. mol.kt t is the